# Dexterous Manipulation through Imitation Learning: A Survey

### v4

@@ III-A1 Description
$\mathcal{L}(\pi)=-\mathbb{E}_{(s,a)\sim p_{\mathcal{D}}}\left[\log\pi(a\mid s)\right]$ (1)
@@ III-A2 Research Progress
BC has achieved significant progress in dexterous manipulation [77, 78, 121, 122, 20] and has demonstrated effective performance in relatively simple tasks, such as pushing [79] and grasping [14]. However, its applicability in dynamic environments and long-horizon tasks remains an active area of research.
@@
The training data for BC models are usually collected from expert demonstrations tailored to specific tasks. Consequently, when the agent encounters states that are unseen during training, it may produce actions that deviate from the expert’s behavior, leading to task failure. In sequential decision-making processes, even small deviations from expert actions at each step can accumulate over time, resulting in what is known as the “compounding error” problem. This issue is particularly pronounced in dexterous manipulation tasks[80, 81], due to the high dimensionality of the action space and the strong dependency between task success and the consistency of the predicted action trajectory. To mitigate compounding errors in dexterous manipulation, Mandlekar et al. [29] propose a hierarchical framework that segments demonstration trajectories at intersection points across different tasks and recombines them to synthesize trajectories for unseen tasks. Similarly, Zhao et al. [77] address the problem by considering the compatibility with high-dimensional visual observations. Instead of predicting actions step by step, they propose to predict entire action sequences, thereby reducing the effective decision horizon and alleviating compounding errors.
@@
Another challenge in BC is its limited ability to model multi-modal data, which is prevalent in human demonstrations collected from real-world environments. To overcome this limitation, several approaches have been proposed to model multi-modal action distributions. Florence et al.[26] formulate BC as a conditional energy-based modeling problem for capturing multi-modal data distribution, albeit at the cost of increased computational overhead. Similarly, Shafiullah et al.[82] propose to model the action distribution as a mixture of Gaussians. Their method leverages the Transformer architecture to efficiently utilize the history of previous observations and enables multi-modal action prediction through token-based outputs. Another promising direction is leveraging generative models to capture the inherent diversity of expert behaviors. Mandlekar et al.[83] propose using generative models for trajectory prediction, enabling selective imitation, though this approach relies on carefully curated, task-specific datasets.
@@
More recently, diffusion models have shown great potential in enhancing the robustness and generalization of BC methods. Chen et al. [84] propose a diffusion-augmented BC framework that models both conditional and joint probability distributions over expert demonstrations. Building on this idea, Chi et al. [32] employ diffusion models as decision models to directly generate sequential actions conditioned on visual input and the robot’s current state. Additionally, the 3D Diffusion Policy [85] leverages 3D input representations to better capture scene spatial configurations. Similarly, the 3D Diffuser Actor [86] utilizes full 3D scene representations by integrating RGB and depth information, along with language instructions, robot proprioception, and noise trajectories, through a 3D Relative Transformer framework.
@@
Compared to variational autoencoders (VAEs) [123] and generative adversarial networks (GANs) [124], diffusion models offer several advantages for imitation learning, particularly in modeling complex, multimodal action distributions [32, 125]. Chen et al. [84] proposed using diffusion models to augment behavioral cloning and conducted a comprehensive comparison with other generative models, including VAEs, GANs, and EBMs(energy-based models)/implicit models, across different implementations in terms of model architecture, sampling/inference cost, and performance on various continuous control and manipulation tasks. Experimental results demonstrated that diffusion-augmented behavioral cloning outperformed standard BC and several other generative model variants on multiple tasks.
@@ III-A3 Discussion
In general, BC-based methods struggle with generalization and modeling multi-modal action distributions. To overcome these limitations, diffusion models have recently attracted increasing attention. They can be employed either as decision models that directly generate action sequences [32] or as high-level strategy models that guide the action generation process [84]. In both settings, diffusion models have shown promising performance and improved flexibility over conventional BC methods.
@@ III-B1 Description
$\psi^{\pi}(s)=\psi_{0}(s)+\gamma\sum_{s^{\prime}\in S}T(s^{\prime}|s,a)\psi^{\pi}(s^{\prime})$ (4)
@@ III-B2 Research Progress
Recent studies have leveraged IRL frameworks to tackle complex dexterous manipulation tasks. Orbik et al. [87] first advanced IRL for dexterous manipulation by introducing reward normalization, task-specific feature masking, and random sample generation. These techniques effectively mitigate reward bias toward demonstrated actions and enhance learning stability in high-dimensional state-action spaces, leading to better generalization across unseen scenarios. Building upon the need for efficient learning in such high-dimensional settings, Generative Causal Imitation Learning [88] improves the sample efficiency of IRL by integrating maximum entropy modeling with adaptive sampling strategies. By leveraging nonlinear function approximation through neural networks, the proposed method enables expressive cost function learning while handling unknown system dynamics. To further incorporate user feedback into the learning process, ErrP-IRL [89] integrates error-related potentials [90] with IRL. This approach assigns trajectory weights based on users’ cognitive responses, which are then used to iteratively refine a reward function represented as a sum of radial basis functions.
Beyond human feedback, recent works have explored learning reward functions from large-scale, unstructured demonstrations. GraphIRL [91] extracts task-specific embeddings from diverse video demonstrations. By modeling object interactions as graphs and performing temporal alignment, GraphIRL learns transferable reward functions without requiring explicit reward design or environment correspondence, enabling cross-domain manipulation capabilities. To further improve policy precision, Naranjo-Campos et al. [92] propose to integrate IRL with Proximal Policy Optimization [52]. Their method incorporates expert-trajectory-based features and a reverse discount factor to address feature vanishing issues near goal states, thereby improving the robustness of the learned policies. More recently, Visual IRL [93] extends the scope of IRL to human-robot collaboration tasks. It employs adversarial IRL to infer reward functions from human demonstration videos and introduces a neuro-symbolic mapping that translates human kinematics into robot joint configurations. This approach not only ensures accurate end-effector placement but also preserves human-like motion dynamics, facilitating natural and effective robot behavior in dexterous manipulation tasks.
@@ III-B3 Discussion
In summary, IRL has demonstrated significant potential for dexterous manipulation tasks. By inferring the underlying reward function from expert demonstrations, IRL enables robots to generalize complex behaviors and adapt to diverse environments without the need for manually designed reward functions. This capability is particularly valuable in dexterous manipulation scenarios where reward specification is challenging or impractical [37, 38]. Despite these promising developments, state-of-the-art IRL methods still face several limitations. One of the primary challenges lies in accurately estimating reward functions, particularly in environments with high-dimensional action spaces or sparse feedback signals. Furthermore, IRL methods often rely on large amounts of expert demonstration data, which poses practical constraints due to the high cost and time required for data collection [37, 94].
@@ III-C1 Description
Generative Adversarial Imitation Learning (GAIL) extends the Generative Adversarial Network (GAN) framework [126] to the domain of IL. It formulates the imitation process as a two-player adversarial game between a generator and a discriminator. The generator corresponds to a policy $\pi$ that aims to produce behavior that closely resembles expert demonstrations, while the discriminator $D(s,a)$ evaluates whether a state-action pair $(s,a)$ originates from the expert data $M$ or is generated by $\pi$.
@@ III-C2 Research Progress
GAIL has been widely adopted in dexterous manipulation. However, its effectiveness heavily relies on the quality and availability of expert demonstrations, which are often labor-intensive to collect and prone to inconsistencies [95, 96]. Such discrepancies arise from factors like collector biases [97], expert errors, noisy data, non-convex solution spaces, and suboptimal strategies [98]. Additionally, data scarcity further limits learning efficiency and policy robustness.
To address these challenges, various GAIL extensions have been proposed. HGAIL [99] employs hindsight experience replay to synthesize expert-like demonstrations without requiring real expert data. AIL-TAC [98] introduces a semi-supervised correction network to refine noisy demonstrations. GAIL has also been used in sim-to-real transfer [100], reducing the dependence on real-world expert data. Nevertheless, GAIL still suffers from mode collapse, where learned policies capture only a narrow range of behaviors, and gradient vanishing issues when the discriminator overpowers the generator. To mitigate these problems, RIDB [101] incorporates variational autoencoders to learn semantic policy embeddings and enable smooth interpolation across behaviors. WAIL [102] leverages the Wasserstein GAN framework [103] to improve training stability and reduce mode collapse. DIL-SOGM [104] further introduces a self-organizing generative model to capture multiple behavioral patterns without requiring encoders. In parallel, several works improve GAIL’s robustness under imperfect demonstrations. GA-GAIL [105] employs a second discriminator to identify goal states, enhancing policy learning from suboptimal data. RB-GAIL [106] integrates ranking mechanisms and multiple discriminators to model diverse behavior modes while leveraging generated experiences.
In addition to addressing the quality and availability of expert demonstrations, recent studies have sought to improve the performance of GAIL-based dexterous manipulation methods in other aspects. For instance, TRAIL [107] introduces constrained discriminator optimization to prevent the discriminator from focusing on spurious, task-irrelevant features such as visual distractors, thereby preserving meaningful reward signals and enhancing task performance. In the context of human imitation, Antotsiou et al. [Antotsiou2018TaskOrientedHM] combine inverse kinematics and particle swarm optimization with GAIL to mitigate sensor noise and domain discrepancies, enabling robots to autonomously grasp objects in simulation environments. Furthermore, P-GAIL [108] incorporates entropy-maximizing deep P-networks into GAIL to improve policy learning in deformable object manipulation tasks.
@@ III-D2 Research Progress
In recent years, HIL has made significant progress in task decomposition and skill generalization. CompILE [109] enhances generalization in complex environments by decomposing tasks into independent sub-tasks, laying the foundation for subsequent work, particularly for tasks with long temporal dependencies. In dexterous manipulation, ARCH [111] integrates a low-level library of predefined skills with a high-level IL policy, enabling efficient skill composition and adaptation for complex, high-precision tasks. Meanwhile, Wan et al. [113] emphasize maintaining skill continuity and stability in dynamic environments by decomposing tasks into continuous sub-tasks and enabling adaptive policy transitions under varying conditions.
To enhance scalability, recent research has increasingly explored automatic hierarchy discovery to reduce reliance on manually designed skill structures. Unsupervised skill segmentation methods have shown strong potential for identifying reusable sub-skills and task boundaries directly from raw data. For example, Gehring et al.[127] propose hierarchical exploration strategies that use unsupervised clustering to learn multi-level reusable sub-policies. More recently, Jansonnie et al.[128] present a skill discovery framework that leverages automatic task generation and asymmetric self-play to produce diverse manipulatory behaviors, which are then embedded into reusable primitives suitable for hierarchical policies. Similarly, contrastive learning approaches [129] discover skills by maximizing mutual information across trajectory embeddings, resulting in more robust and generalizable skill representations.
From a data-centric perspective, play data has been utilized to train both high-level and low-level policies. Wang et al.[114] propose MimicPlay, which exploits unstructured human play interactions to learn high-level latent plans and trains a low-level visuomotor controller using only a small number of demonstrations. Likewise, Lin et al.[115] introduce H2RIL, which extracts interaction-aware skill embeddings from task-agnostic play data and aligns them with human videos via temporal contrastive learning, enabling generalization to novel, composable tasks and adaptation to out-of-distribution scenarios.s
@@ III-D3 Discussion
In summary, HIL has demonstrated significant advantages in task decomposition, skill generalization, and handling long-horizon tasks. By introducing hierarchical structures and multi-level control strategies, substantial progress has been made across various dexterous manipulation tasks. However, most current approaches still depend on manually designed hierarchies, which limits autonomy in unstructured or dynamic environments. Automatic hierarchy discovery methods, such as unsupervised skill segmentation or contrastive learning-based sub-task discovery, are essential for enabling robots to autonomously identify reusable skills and adapt their hierarchy to new tasks. Furthermore, ensuring robustness and continuity when skill libraries need to be updated or extended, particularly under changing task conditions, remains an open problem.
@@ III-E2 Research Progress
Various CIL approaches have been developed for dexterous manipulation, each aiming to balance memory retention and adaptation. Early studies [116] relied on storing large amounts of task data, leading to high storage and computational costs. To address this, task-specific adapter structures introduced lightweight, modular components for seamless task switching [117], though their performance declines when task variations are substantial.
To address catastrophic forgetting more systematically, regularization-based methods[130] have gained prominence. For instance, Elastic Weight Consolidation (EWC)[131] penalizes large updates to parameters critical for previously learned tasks by incorporating Fisher information into the loss function. Similarly, knowledge distillation-based approaches[132] preserve prior skills by encouraging new policies to mimic outputs from older policies. Rehearsal-based strategies offer another solution, such as experience replay, where past experiences or generated trajectories are interleaved with new task data to reduce forgetting[133]. Generative models, such as Deep Generative Replay (DGR), further alleviate storage constraints by synthesizing realistic task trajectories instead of storing raw data [119], although ensuring trajectory fidelity remains challenging.
Beyond these, unsupervised skill discovery has been explored to improve adaptability and generalization [113]. It dynamically generates new skills and integrates them into the robot’s repertoire, but has yet to demonstrate robust real-world generalization. Unified policy learning via behavior distillation [118] provides an alternative by employing a single shared policy across tasks, eliminating the need for task-specific modules, though at the cost of increased task interference. Lastly, self-supervised learning techniques have shown promise in acquiring transferable skill abstractions without explicit demonstrations  [120].

### v5

@@ III-A1 Description
$\mathcal{L}(\pi)=-\mathbb{E}_{(s,a)\sim p_{\mathcal{D}}}\left[\log\pi(a\mid s)\right].$ (1)
@@ III-A2 Research Progress
BC has achieved significant progress in dexterous manipulation [zhao2023learning, arunachalam2023dexterous, 10064325, 10802784, qin2022one] and has demonstrated effective performance in relatively simple tasks, such as pushing [8462901] and grasping [zhang2018deep]. However, its applicability in dynamic environments and long-horizon tasks remains an active area of research.
@@
The training data for BC models are usually collected from expert demonstrations tailored to specific tasks. Consequently, when the agent encounters states that are unseen during training, it may produce actions that deviate from the expert’s behavior, leading to task failure. In sequential decision-making processes, even small deviations from expert actions at each step can accumulate over time, resulting in what is known as the “compounding error” problem. This issue is particularly pronounced in dexterous manipulation tasks[osa2018algorithmic, ke2021grasping], due to the high dimensionality of the action space and the strong dependency between task success and the consistency of the predicted action trajectory. To mitigate compounding errors in dexterous manipulation, Mandlekar et al. [mandlekar2020learning] proposed a hierarchical framework that segments demonstration trajectories at intersection points across different tasks and recombines them to synthesize trajectories for unseen tasks. Similarly, Zhao et al. [zhao2023learning] addressed the problem by considering the compatibility with high-dimensional visual observations. Instead of predicting actions step by step, they propose to predict entire action sequences, thereby reducing the effective decision horizon and alleviating compounding errors.
@@
Another challenge in BC is its limited ability to model multimodal data, which is prevalent in human demonstrations collected from real-world environments. To overcome this limitation, several approaches have been proposed to model multi-modal action distributions. Florence et al.[florence2022implicit] formulated BC as a conditional energy-based modeling problem for capturing multi-modal data distribution, albeit at the cost of increased computational overhead. Similarly, Shafiullah et al.[shafiullah2022behavior] proposed to model the action distribution as a mixture of Gaussians. Their method leverages the Transformer architecture to efficiently utilize the history of previous observations and enables multi-modal action prediction through token-based outputs. Another promising direction is leveraging generative models to capture the inherent diversity of expert behaviors. Mandlekar et al.[9196935] proposed using generative models for trajectory prediction, enabling selective imitation, though this approach relies on carefully curated, task-specific datasets.
@@
More recently, diffusion models have shown great potential in enhancing the robustness and generalization of BC methods. Chen et al. [chen2023diffusion] proposed a diffusion-augmented BC framework that models both conditional and joint probability distributions over expert demonstrations. Building on this idea, Chi et al. [chi2023diffusion] employed diffusion models as decision models to directly generate sequential actions conditioned on visual input and the robot’s current state. Additionally, the 3D Diffusion Policy [ze2024d] leveraged 3D input representations to better capture scene spatial configurations. Similarly, the 3D Diffuser Actor [ke20243d] utilized full 3D scene representations by integrating RGB and depth information, along with language instructions, robot proprioception, and noise trajectories, through a 3D Relative Transformer framework.
@@
Compared to variational autoencoders (VAEs) [kingma2019introduction] and generative adversarial networks (GANs) [goodfellow2020generative], diffusion models offer several advantages for imitation learning, particularly in modeling complex, multimodal action distributions [chi2023diffusion, yan2025m]. Chen et al. [chen2023diffusion] proposed using diffusion models to augment BC and conducted a comprehensive comparison with other generative models, including VAEs, GANs, and EBMs (energy-based models)/implicit models, across different implementations in terms of model architecture, sampling/inference cost, and performance on various continuous control and manipulation tasks. Experimental results demonstrated that diffusion-augmented BC outperformed standard BC and several other generative model variants on multiple tasks.
@@ III-A3 Discussion
In general, BC-based methods struggle with generalization and modeling multi-modal action distributions. To overcome these limitations, diffusion models have recently attracted increasing attention. They can be employed either as decision models that directly generate action sequences [chi2023diffusion] or as high-level strategy models that guide the action generation process [chen2023diffusion]. In both settings, diffusion models have shown promising performance and improved flexibility over conventional BC methods.
@@ III-B1 Description
$\psi^{\pi}(s)=\psi_{0}(s)+\gamma\sum_{s^{\prime}\in S}T(s^{\prime}|s,a)\psi^{\pi}(s^{\prime}).$ (4)
@@ III-B2 Research Progress
Recent studies have leveraged IRL frameworks to tackle complex dexterous manipulation tasks. Orbik et al. [9515637] first advanced IRL for dexterous manipulation by introducing reward normalization, task-specific feature masking, and random sample generation. These techniques effectively mitigate reward bias toward demonstrated actions and enhance learning stability in high-dimensional state-action spaces, leading to better generalization across unseen scenarios. Building upon the need for efficient learning in such high-dimensional settings, Generative Causal Imitation Learning [1603.00448] improved the sample efficiency of IRL by integrating maximum entropy modeling with adaptive sampling strategies. By leveraging nonlinear function approximation through neural networks, the proposed method enables expressive cost function learning while handling unknown system dynamics. To further incorporate user feedback into the learning process, ErrP-IRL [batzianoulis2021] integrated error-related potentials [millan2008] with IRL. This approach assigns trajectory weights based on users’ cognitive responses, which are then used to iteratively refine a reward function represented as a sum of radial basis functions.
Beyond human feedback, recent works have explored learning reward functions from large-scale, unstructured demonstrations. GraphIRL [2207.14299] extracted task-specific embeddings from diverse video demonstrations. By modeling object interactions as graphs and performing temporal alignment, GraphIRL learns transferable reward functions without requiring explicit reward design or environment correspondence, enabling cross-domain manipulation capabilities. To further improve policy precision, Naranjo-Campos et al. [app142311131] proposed to integrate IRL with Proximal Policy Optimization [schulman2017proximal]. Their method incorporates expert-trajectory-based features and a reverse discount factor to address feature vanishing issues near goal states, thereby improving the robustness of the learned policies. More recently, Visual IRL [2412.11360] extended the scope of IRL to human-robot collaboration tasks. It employs adversarial IRL to infer reward functions from human demonstration videos and introduces a neuro-symbolic mapping that translates human kinematics into robot joint configurations. This approach not only ensures accurate end-effector placement but also preserves human-like motion dynamics, facilitating natural and effective robot behavior in dexterous manipulation tasks.
@@ III-B3 Discussion
In summary, IRL has demonstrated significant potential for dexterous manipulation tasks. By inferring the underlying reward function from expert demonstrations, IRL enables robots to generalize complex behaviors and adapt to diverse environments without the need for manually designed reward functions. This capability is particularly valuable in dexterous manipulation scenarios where reward specification is challenging or impractical [ARORA2021103500, s23073762]. Despite these promising developments, state-of-the-art IRL methods still face several limitations. One of the primary challenges lies in accurately estimating reward functions, particularly in environments with high-dimensional action spaces or sparse feedback signals. Furthermore, IRL methods often rely on large amounts of expert demonstration data, which poses practical constraints due to the high cost and time required for data collection [ARORA2021103500, 10493015].
@@ III-C1 Description
Generative Adversarial Imitation Learning (GAIL) extends the Generative Adversarial Network (GAN) framework [goodfellow2014generative] to the domain of IL. It formulates the imitation process as a two-player adversarial game between a generator and a discriminator. The generator corresponds to a policy $\pi$ that aims to produce behavior that closely resembles expert demonstrations, while the discriminator $D(s,a)$ evaluates whether a state-action pair $(s,a)$ originates from the expert data $M$ or is generated by $\pi$.
@@ III-C2 Research Progress
GAIL has been widely adopted in dexterous manipulation. However, its effectiveness heavily relies on the quality and availability of expert demonstrations, which are often labor-intensive to collect and prone to inconsistencies [tangkaratt2019vild, wang2021learning]. Such discrepancies arise from factors like collector biases [zuo2021adversarial], expert errors, noisy data, non-convex solution spaces, and suboptimal strategies [antotsiou2021adversarial]. Additionally, data scarcity further limits learning efficiency and policy robustness.
To address these challenges, various GAIL extensions have been proposed. HGAIL [liu2019hindsight] employed hindsight experience replay to synthesize expert-like demonstrations without requiring real expert data. AIL-TAC [antotsiou2021adversarial] introduced a semi-supervised correction network to refine noisy demonstrations. GAIL has also been used in sim-to-real transfer [jiang2023mastering], reducing the dependence on real-world expert data. Nevertheless, GAIL still suffers from mode collapse, where learned policies capture only a narrow range of behaviors, and gradient vanishing issues when the discriminator overpowers the generator. To mitigate these problems, RIDB [wang2017robust] incorporated variational autoencoders to learn semantic policy embeddings and enable smooth interpolation across behaviors. WAIL [xiao2019wasserstein] leveraged the Wasserstein GAN framework [arjovsky2017wasserstein] to improve training stability and reduce mode collapse. DIL-SOGM [vahabpour2024diverse] further introduced a self-organizing generative model to capture multiple behavioral patterns without requiring encoders. In parallel, several works improve GAIL’s robustness under imperfect demonstrations. GA-GAIL [tsurumine2022goal] employed a second discriminator to identify goal states, enhancing policy learning from suboptimal data. RB-GAIL [shi2024ranking] integrated ranking mechanisms and multiple discriminators to model diverse behavior modes while leveraging generated experiences.
In addition to addressing the quality and availability of expert demonstrations, recent studies have sought to improve the performance of GAIL-based dexterous manipulation methods in other aspects. For instance, TRAIL [zolna2021task] introduced constrained discriminator optimization to prevent the discriminator from focusing on spurious, task-irrelevant features such as visual distractors, thereby preserving meaningful reward signals and enhancing task performance. In the context of human imitation, Antotsiou et al. [antotsiou2018task] combined inverse kinematics and particle swarm optimization with GAIL to mitigate sensor noise and domain discrepancies, enabling robots to autonomously grasp objects in simulation environments. Furthermore, P-GAIL [tsurumine2019generative] incorporated entropy-maximizing deep P-networks into GAIL to improve policy learning in deformable object manipulation tasks.
@@ III-D2 Research Progress
In recent years, HIL has made significant progress in task decomposition and skill generalization. CompILE [Kipf2018CompositionalIL] enhanced generalization in complex environments by decomposing tasks into independent sub-tasks, laying the foundation for subsequent work, particularly for tasks with long temporal dependencies. In dexterous manipulation, ARCH [2409.16451] integrated a low-level library of predefined skills with a high-level IL policy, enabling efficient skill composition and adaptation for complex, high-precision tasks. Meanwhile, Wan et al. [Wan2023LOTUSCI] emphasized maintaining skill continuity and stability in dynamic environments by decomposing tasks into continuous sub-tasks and enabling adaptive policy transitions under varying conditions.
To enhance scalability, recent research has increasingly explored automatic hierarchy discovery to reduce reliance on manually designed skill structures. Unsupervised skill segmentation methods have shown strong potential for identifying reusable sub-skills and task boundaries directly from raw data. For example, Gehring et al.[gehring2021hierarchical] proposed hierarchical exploration strategies that use unsupervised clustering to learn multi-level reusable sub-policies. More recently, Jansonnie et al.[jansonnie2024unsupervised] presented a skill discovery framework that leverages automatic task generation and asymmetric self-play to produce diverse manipulatory behaviors, which are then embedded into reusable primitives suitable for hierarchical policies. Similarly, contrastive learning approaches [yang2023behavior] discovered skills by maximizing mutual information across trajectory embeddings, resulting in more robust and generalizable skill representations.
From a data-centric perspective, play data has been used to train both high-level and low-level policies. Wang et al.[wang2023mimicplay] proposed MimicPlay, which exploits unstructured human play interactions to learn high-level latent plans and trains a low-level visuomotor controller using only a small number of demonstrations. Lin et al.[10610084] introduced H2RIL, which extracts interaction-aware skill embeddings from task-agnostic play data and aligns them with human videos via temporal contrastive learning, enabling generalization to novel, composable tasks and adaptation to out-of-distribution scenarios.
@@ III-D3 Discussion
In summary, HIL has demonstrated significant advantages in task decomposition, skill generalization, and handling long-horizon tasks. By introducing hierarchical structures and multi-level control strategies, substantial progress has been made across various dexterous manipulation tasks. However, most current approaches still depend on manually designed hierarchies, which limit autonomy in unstructured or dynamic environments. Automatic hierarchy discovery methods, such as unsupervised skill segmentation or contrastive learning-based sub-task discovery, are essential for enabling robots to autonomously identify reusable skills and adapt their hierarchy to new tasks. Furthermore, ensuring robustness and continuity when skill libraries need to be updated or extended, particularly under changing task conditions, remains an open problem.
@@ III-E2 Research Progress
Various CIL approaches have been developed for dexterous manipulation, each aiming to balance memory retention and adaptation. Early studies [Liang2024NeverEndingBA] relied on storing large amounts of task data, leading to high storage and computational costs. To address this, task-specific adapter structures introduced lightweight, modular components for seamless task switching [Liu2023TAILTA], though their performance declined when task variations are substantial.
To address catastrophic forgetting more systematically, regularization-based methods[zhaostatistical] have gained prominence. For instance, Elastic Weight Consolidation (EWC)[kirkpatrick2017overcoming] penalized large updates to parameters critical for previously learned tasks by incorporating Fisher information into the loss function. Similarly, knowledge distillation-based approaches[li2024continual] preserved prior skills by encouraging new policies to mimic outputs from older policies. Rehearsal-based strategies offered another solution, such as experience replay, where past experiences or generated trajectories are interleaved with new task data to reduce forgetting[rolnick2019experience]. Generative models, such as Deep Generative Replay (DGR), further alleviated storage constraints by synthesizing realistic task trajectories instead of storing raw data [Gao2021CRILCR], although ensuring trajectory fidelity remains challenging.
Beyond these, unsupervised skill discovery has been explored to improve adaptability and generalization [Wan2023LOTUSCI]. It dynamically generated new skills and integrated them into the robot’s repertoire, but had yet to demonstrate robust real-world generalization. Unified policy learning via behavior distillation [Haldar2023PolyTaskLU] provided an alternative by employing a single shared policy across tasks, eliminating the need for task-specific modules, though at the cost of increased task interference. Lastly, self-supervised learning techniques have shown promise in acquiring transferable skill abstractions without explicit demonstrations  [Mete2024QueSTSS].